OpenStudy (anonymous):

In an atom two electrons move round the nucleus in circular orbits of radii R and 4R respectively. The ratio of the time taken by them to complete one revolution is

OpenStudy (espex):

If the one orbit is 4 times the length of the first, it seems that it would be a 1:4 ratio.

OpenStudy (anonymous):

it is 1:8

OpenStudy (espex):

Because it is a radius..

OpenStudy (hoblos):

you must find the ratio of the parameters

OpenStudy (anonymous):

t^2 ~r^3

OpenStudy (espex):

hoblos: Could you elaborate please.

OpenStudy (anonymous):

square of the time period directly proportional to the cube of the radius

OpenStudy (espex):

That is for planetary orbits which electrons do not mimic. I was under the impression it was a probability function that defined electron orbits.

OpenStudy (anonymous):

Assuming that they move with the same speed then hoblos is right because the time taken for 1 revolution is directly porportional to the perimeter of the circular paths. therefore the perimeter of the electron with radii R is \[2\pi r\] and the perimeter of the other electron is \[2\pi(4R)=8\pi r\] Therefore, \[8\pi r:2\pi r=4\]